# RacerF: Lightweight Static Data Race Detection for C Code

### v1

@@ 5 Thread Analysis
In this section, we propose our approach how to leverage an analyser for sequential programs to analyse programs with multiple threads. The design of our approach is based on the observation that answers to our queries described in Section 4 on objects relevant for our analysis are often not too much affected by thread interleavings. We design two strategies of combining results of the analyses of the sequential behaviour of individual threads that under- and over-approximate the overall behaviour of the concurrent program consisting of the threads. Furthermore, these strategies can subsequently be combined to increase precision of the resulting analysis by using under-approximation to find races while using over-approximation to show their absence. However, as we show in our experimental evaluation in Section 8, both of our strategies are often sufficiently precise on their own.
The core idea of our thread analysis is similar to that of the Frama-C’s proprietary Mthread plugin described in [28]. Starting with the $\mathtt{main}$ thread, we iteratively build a thread-creation graph that captures the parent-child relationship among the so-far discovered threads. In each iteration, we add newly discovered threads and perform the analysis (potentially multiple times) of each thread as a sequential program (using the $\mathtt{analyse\_thr}$ function of the selected backend). This is done by solving a system of equations over the initial states of the threads given by whose construction we describe below. However, before that, we first describe the graph construction algorithm.
@@
$R_{0}=\{\mathsf{main}\mapsto\mathsf{Backend.analyse\_thr}(\mathsf{main},\mathsf{GlobalsVals})\}$
@@ 18       $I_{i}=\mathsf{SolveEquations}(\Phi_{i},I_{i-1})$
20       $R_{i}=\{t\mapsto\mathsf{Backend.analyse}(t,I_{i}(t))\;|\;t\in\mathsf{dom}(I_{i})\}$
@@ 5.1 Construction of the Thread-Creation Graph
After a new iteration of the graph $G_{i}$ is constructed, we build a system of equations, one for each discovered thread, given by $G_{i}$. Each of these equations describes how an initial state of a thread can be computed using initial states of its parents. Then, the system is solved using initial states from the previous iteration $I_{i-1}$ to obtain new initial states $I_{i}$. The solution may need application of widening operators – we provide more details on our implementation of the equation solving in Section 8. Based on the new initial states, we compute new results $R_{i}$ on line 14.
This loop is repeated until the thread-creation graph stabilises. Overall, the algorithm runs two fixpoint loops – the outer loop updates the graph with new threads, while the inner loop performs analysis on the current graph, and can thus help to discover previously unknown threads in new iterations.
@@ 5.2 Under-Approximating Thread Analysis
We use the following under-approximated semantics of thread creation. A thread is started with the initial state corresponding to the join of initial states of its create statements, and its writes to global variables are not propagated to other threads. The creation of threads then roughly corresponds to calling its entry point in a “sandbox” that makes all of its actions invisible for other threads. Formally, given a thread graph $G_{i}$ with edges $\xrightarrow{}$, we compute the initial state of a thread $t$ (other than $\mathtt{main}$ – the initial state of $\mathtt{main}$ is given by the initial states of global variables) as follows:
@@
The motivation for this strategy is straightforward – for acyclic thread-creation graphs, it is enough to analyse each entry point only once (provided caching is used for the $\mathtt{analyse\_thr}$ function). On the other hand, there are two main sources of possible imprecisions – thread interleavings are not taken into account, and parents do not see effects of their children. The latter can sometimes result in seemingly weird false negatives when performing data race detection, as demonstrated in Listing 3(a). In the program, the $\mathtt{main}$ thread first creates a thread $\mathtt{t1}$ which allocates an integer value and then a thread $\mathtt{t2}$ (created twice) accesses this value without any synchronization. However, we cannot detect this data race using EVA [^2] We could detect the memory access with the other backends that do not care about memory initialization. since in our under-approximated setting, $\mathtt{t2}$ will never see the allocation, and in such a situation, EVA will mark the memory access on line 6 as surely invalid and will not report any memory access for it.
@@
[linenos]c int *g; void *t1() g = malloc(…); void *t2() (*g)++; int main() create(t1); join(t1); create(t2); create(t2);
@@
[linenos]c int g = 1; void *thread() lock(L); if (g == 1) unlock(L); return NULL; g++; int main() create(t1); lock(L); g = 0; g = 1; unlock(L);
@@ 5.3 Over-Approximating Thread Analysis
We compute the initial state of a thread $t$ (here also including the $\mathtt{main}$ thread) as the join of the states our analysis has encountered so far. Formally, using the set $T$ of currently known threads and the initial values of global variables $V=\mathsf{GlobalsVals}$, we proceed as follows:
$I(t)\triangleq V\;\sqcup\bigsqcup_{s\in\mathsf{Stmt},\;t^{\prime}\in T}\mathsf{let}\;R=\mathsf{Backend.analyse\_thr}(t^{\prime},I(t^{\prime}))\;\mathsf{in}\;\mathsf{Backend.state}(R,(t^{\prime},\varepsilon,s)),$
@@
As we also consider context switches that cannot happen due to synchronization (even those that cannot happen because a child was not yet created, e.g., when computing the initial state of the main thread and already using states of its children), we may later report false data races as demonstrated in Listing 3(b). In this example, we will compute the initial state of $\mathtt{thread}$ to be $\{\mathtt{g}\mapsto\{0,1\}\}$. However, the value 00 is never visible to $\mathtt{thread}$ because the sequence of updates on line 10 (as well as the condition on line 4) is performed atomically, resulting in no change of $\mathtt{g}$. Thus, the memory access on line 6, which would be flagged as a data race (together with one of the accesses on line 10), is in fact unreachable.
@@
[numbersep=-6pt, obeytabs=true,linenos, tabsize=1]c atomic_int g; void *t(…) g++; int main() g = 0; create(t); if (g == 1) …
Figure 4: An example of a program for which the over-approximating strategy fails to compute a sound over-approximation unless the program is transformed.
@@
Program transformation.
The over-approximating strategy is not in itself sufficient to compute a sound over-approximation of multi-threaded behaviour as demonstrated on the program in Figure 4. In the example, we are particularly interested in the block inside the condition on line 11, which may contain some important event for our analyser. This code is unreachable when the main thread is analysed as a sequential program regardless of the initial state. Indeed, our over-approximation strategy would compute the initial state as $\{\mathtt{g}\mapsto[0,\infty]\}$, but already on line 6, this state would be modified to $\{\mathtt{g}\mapsto[0,0]\}$ and the over-approximation would be lost.
@@
  • An assignment $\mathtt{x\coloneq y}$ is transformed into a non-deterministic choice between the old and the new value as $\mathtt{x\coloneq(\ast)\;?\;x:y}$.
@@ 6.1 Lockset Analysis
Lockset analysis is traditionally used to compute must- or may-locksets being held for each statement. Our lockset analysis is context-sensitive, path-insensitive, and it computes both must- and may-locksets at once. It is inspired by the approach of RacerX described in [13], which, however, does not employ any advanced alias analysis and is not context-sensitive.
We represent locks by their concrete addresses, i.e., $\mathsf{Lock}=\mathsf{Address}$, and we define $\mathsf{Lockset}=2^{\mathsf{Lock}}$. Our lockset analysis is implemented as a forward data flow analysis with the domain of locksets, computed for each thread in separation, starting with the empty lockset. On the intraprocedural level, we start to analyse a function with an entry lockset $L_{entry}$ coming from the caller, and gradually update it using the following transfer function that, for a context $c$ with a statement $s$ and using $P=\mathsf{Backend.eval\_ptr}(c,p)$, transforms the lockset $L_{old}$ reached before executing $s$ into a set of locksets $\mathcal{L}_{new}$ resulting from the execution of $s$ as follows: $\mathcal{L}_{new}=L_{old}\cup\{\ell\}\;|\;\ell\in P\}$ if $s$ is $\mathtt{lock}(p)$, $\mathcal{L}_{new}=\{L_{old}\setminus\,\{\ell\}\;|\;\ell\in P\}$ if $s$ is $\mathtt{unlock}(p)$, and $\mathcal{L}_{new}=\{L_{old}\}$ otherwise.
After applying the transfer function, if $\mathcal{L}_{\mathrm{new}}$ is not a singleton set, we fork the analysis and continue with each element of $\mathcal{L}_{\mathrm{new}}$ separately. As a result, we obtain the summary of a function $f$ analysed in a context $c$ in the form $(c,L_{entry})\mapsto\mathcal{L}_{exit}$ where $\mathcal{L}_{exit}$ is the set of locksets at the return statements of all forked analysis runs inside $f$. Such summaries are then used for the interprocedural analysis. We also construct statement summaries $\mathcal{S}$ of the same form for individual statements. These are used for caching at the statement level and to implement the must- and may-lockset queries for a context $c$ as $\mathsf{may\_ls}(c)=\bigcup_{(c,L)\in\mathsf{dom}(\mathcal{S})}L$ and $\mathsf{must\_ls}(c)=\bigcap_{(c,L)\in\mathsf{dom}(\mathcal{S})}L$.
@@
As we represent locks using memory addresses and fork the analysis whenever we cannot precisely evaluate the parameter of a (un)locking operation, the set of traversed paths can easily explode. To tackle this problem, our transfer function selects only a small number of locks when the set $\mathsf{Backend.eval\_ptr}(c,p)$ is too large (currently, the threshold is set to three locks). This works in many practical cases because it is usually enough to differentiate the may- and must-cases through a few paths only. In the future, we would like to provide support for symbolic treatment of locks combined with may- and must-equality reasoning.
@@
For non-blocking lock functions such as $\mathtt{trylock}$, which does not block when mutex is already locked, or $\mathtt{timedlock}$, which blocks just for a limited time, we update the lockset with relevant locks for which we also track the variable used to store the return value of the corresponding call. Evaluation of a query $\mathsf{may\_ls}(c)$ or $\mathsf{must\_ls}(c)$ will then ask the backend for the state $s$ computed for the context $c$ through the query $s=\mathsf{Backend.state}(c)$, and filter out the locks for which the state $s$ does not guarantee that the locking was (possibly or surely, depending on the query) successful. This way we obtain a limited path-sensitivity for branching over expressions containing variables storing return values of lock operations. For the function $\mathtt{pthread\_mutex\_lock}$, we assume that it is blocking (i.e., that it never fails), but the user can easily change the configuration to treat it as non-blocking.
@@
To support read-write locks, a lock object keeps additional information about whether it was obtained in the read- or write-mode. Based on this information, we modify the intersection of locksets, presented later on as the main operation used for data race detection, so that it contains only those locks which are present in both operands and at least in one of them in the write-mode.
@@ 6.2 Active-Threads Analysis
The sets of active threads are computed in a very similar way to how the lockset analysis is performed – just working with sets of threads instead of sets of locks, with the transfer functions for thread creation and joining analogical to those for locking/unlocking, and with the same way of handling the may- and must-queries. The set of thread entry points is usually small, and we thus do not need any special treatment to tackle path explosion unlike in the lockset analysis. One difference between the analyses is in the initial states. The lockset analysis of each thread is started with the empty lockset, but the active-threads analysis of a thread $t$ starts with the union of all threads that may be active together with $t$ in some context of the parent of $t$ (the analysis is thus run in the order given by a topological order on thread-creation relation, and for the very rare case of cyclic thread-creation graphs, we simply skip the analysis, providing trivial answers for queries based on it).
@@ 7 Data Race Detection
An important part of data race detection is discovering local bases that may escape their scope. Our syntactic backend assumes that dynamically allocated bases can always escape and static bases escape when their address is taken. With the Alias and EVA backends, we improve this by testing not only which addresses were taken but also assigned to global variables. (also for dynamic bases).
May and must data races.
We distinguish two types of data races based on the confidence we have in them. For may data races, the below conditions need to be satisfied:  (1) At least one of the accesses is a write access. (2) The accessing threads are either distinct or the same non-unique thread. (3) The offsets of the addresses may overlap (checked by a query over the $\mathsf{CValue}$ domain of intervals). (4) The intersection of the may-locksets of the concerned contexts is empty (checked using the results of the lockset analysis). (5) The threads of both accesses may run in parallel (checked using the results of the active-threads analysis).
For must data races, we require the must-variants of Conditions 3–6 to hold. In addition, the following conditions are required to rule out races involving a memory access representing abstraction of multiple memory accesses:  (6) The memory base is not weak (i.e., if it is dynamic, then it represents precisely one dynamic allocation) and it is not allocated in a non-unique thread. (7) The size of the access does not exceed the size of the type of the base. (8) The base is the only base being read/written to in its context. Condition 6 ensures that we do not report must races on bases possibly obtained by different allocations. Condition 7 rules out races on possibly different indices to arrays. Finally, Condition 8 handles situations when our context-sensitivity level is not sufficient to distinguish accesses to multiple bases.

### v2

@@ 5 Thread Analysis
In this section, we propose our approach on how to leverage an analyser for sequential programs to analyse programs with multiple threads. The design of our approach is based on the observation that answers to our queries described in Section 4 on objects relevant for our analysis are often not too much affected by thread interleavings. We design two strategies for combining results of the analyses of the sequential behaviour of individual threads that under- and over-approximate the overall behaviour of the concurrent program consisting of the threads (not including behaviour possibly caused by errors in the code, i.e., for example undefined behaviour after a data race). Furthermore, these strategies can subsequently be combined to increase the precision of the resulting analysis by using under-approximation to find races while using over-approximation to show their absence. However, as we show in our experimental evaluation in Section 8, both of our strategies are often sufficiently precise on their own.
The core idea of our thread analysis is similar to that of the Frama-C’s proprietary Mthread plugin described in [33]. Starting with the $\mathtt{main}$ thread, we iteratively build a thread-creation graph that captures the parent-child relationship among the so-far discovered threads. In each iteration, we add newly discovered threads and perform the analysis (potentially multiple times) of each thread as a sequential program (using the $\mathtt{analyse\_thread}$ function of the selected backend). This is done by solving a system of equations over the initial states of the threads given by whose construction we describe below. However, before that, we first describe the graph construction algorithm.
@@
$R_{0}=\{\mathsf{main}\mapsto\mathsf{Backend.analyse\_thread}(\mathsf{main},\mathsf{GlobalsVals})\}$
@@ 18       $I_{i}=\mathsf{SolveEquations}(\Phi_{i},I_{i-1})$
20       $R_{i}=\{t\mapsto\mathsf{Backend.analyse\_thread}(t,I_{i}(t))\;|\;t\in\mathsf{dom}(I_{i})\}$
@@ 5.1 Construction of the Thread-Creation Graph
After a new iteration of the graph $G_{i}$ is constructed, we build a system of equations, one for each discovered thread, given by $G_{i}$. Each of these equations describes how an initial state of a thread can be computed using initial states of other threads. Then, the system is solved using initial states from the previous iteration $I_{i-1}$ to obtain new initial states $I_{i}$. The solution may need an application of widening operators – we provide more details on our implementation of the equation solving in Section 8. Based on the new initial states, we compute new results $R_{i}$ on line 14.
This loop is repeated until the thread-creation graph stabilises. Overall, the algorithm runs two fixpoint computations – the outer updates the graph with new threads, while the inner (equation solving on line 13) performs analysis on the current graph, and can thus help to discover previously unknown threads in new iterations.
@@ 5.2 Under-Approximating Thread Analysis
We use the following under-approximated semantics of thread creation. A thread is started with the initial state corresponding to the join of initial states of its create statements, and its writes to global variables are not propagated to other threads. The creation of threads then roughly corresponds to calling its entry point in a “sandbox” that makes all of its actions invisible for other threads. Formally, given a thread-create graph $G_{i}$ with edges $\xrightarrow{}$, we compute the initial state of a thread $t$ (other than $\mathtt{main}$ – the initial state of $\mathtt{main}$ is given by the initial states of global variables) as follows:
@@
The motivation for this strategy is straightforward – for acyclic thread-creation graphs, it is enough to analyse each entry point only once (provided caching is used for the $\mathtt{analyse\_thread}$ function). On the other hand, there are two main sources of possible imprecisions – thread interleavings are not taken into account, and parents do not see effects of their children. The latter can sometimes result in seemingly weird false negatives when performing data race detection, as demonstrated in Listing 4(a). In the program, the $\mathtt{main}$ thread first creates a thread $\mathtt{t1}$ which allocates an integer value and then a thread $\mathtt{t2}$ (created twice) accesses this value without any synchronization. However, we cannot detect this data race using EVA [^2] We could detect the memory access with the other backends that do not care about memory initialization. since in our under-approximated setting, $\mathtt{t2}$ will never see the allocation, and in such a situation, EVA will mark the memory access on line 6 as surely invalid and will not report any memory access for it.
@@
[numbersep=-6pt, obeytabs=true, linenos=true]c int *g; void *t1() g = malloc(…); void *t2() (*g)++; int main() create(t1); join(t1); create(t2); create(t2);
@@
[numbersep=-6pt, obeytabs=true, linenos]c int g = 1; void *thread() lock(L); if (g == 1) unlock(L); return NULL; g = 2; int main() create(t1); lock(L); g = 0; g = 1; unlock(L);
@@ 5.3 Over-Approximating Thread Analysis
We compute the initial state of a thread $t$ (here also including the $\mathtt{main}$ thread) as the join of the states our analysis has encountered so far. Formally, given a thread graph $G_{i}$ with a set $T$ of currently known threads, the initial values of global variables $V=\mathsf{GlobalsVals}$, and a set of all program statements $\mathsf{Stmt}$, we proceed as follows:
$I(t)\triangleq V\;\sqcup\bigsqcup_{s\in\mathsf{Stmt},\;t^{\prime}\in T}\mathsf{let}\;R=\mathsf{Backend.analyse\_thread}(t^{\prime},I(t^{\prime}))\;\mathsf{in}\;\mathsf{Backend.state}(R,(t^{\prime},\varepsilon,s)),$
@@
As we also consider context switches that cannot happen due to synchronization (even those that cannot happen because a child was not yet created, e.g., when computing the initial state of the main thread and already using states of its children), we may later report false data races as demonstrated in Listing 4(b). In this example, we will compute the initial state of $\mathtt{thread}$ to be $\{\mathtt{g}\mapsto\{0,1,2\}\}$. However, the value 00 is never visible to $\mathtt{thread}$ because the sequence of updates on line 10 (as well as the condition on line 4) is performed atomically, resulting in no change of $\mathtt{g}$. Thus, the memory access on line 6, which would be flagged as a data race (together with one of the accesses on line 10), is in fact unreachable.
@@
[numbersep=-6pt, obeytabs=true,linenos, tabsize=1]c atomic_int g;
void *t() g++;
int main() g = 0; create(t); if (g == 1) …
(a) Original program.
{minted}
[obeytabs=true, linenos=false, tabsize=1, highlightlines=4,8, highlightcolor=lightyellow]c atomic_int g;
void *t() g = (*) ? g : g + 1;
int main() g = (*) ? g : 0; create(t); if (g == 1) …
(b) Transformed program.
[…]
@@
5.3.1 Program transformation
The presented over-approximating strategy is not in itself sufficient to compute a sound over-approximation of multi-threaded behaviour as demonstrated on the program in Figure 5. In the example, we are particularly interested in the block inside the condition on line 10, which may contain some important event for our analyser. This code is unreachable when the main thread is analysed as a sequential program regardless of the initial state. Indeed, our over-approximation strategy would compute the initial state as $\{\mathtt{g}\mapsto[0,\mathsf{max\_int}]\}$, but already on line 8, this state would be modified to $\{\mathtt{g}\mapsto[0,0]\}$ and the over-approximation would be lost.
@@ 5.3.1 Program transformation
  • An assignment $\mathtt{x=y}$ is transformed into a non-deterministic choice between the old and the new value as $\mathtt{x=(\ast)\;?\;x:y}$.
@@ 6.1 Lockset Analysis
Lockset analysis is traditionally used to compute must- or may-locksets being held for each statement. Our lockset analysis is context-sensitive, path-insensitive, and it computes both must- and may-locksets simultaneously. It is inspired by the approach of RacerX described in [14], which, however, does not employ any advanced alias analysis and is not context-sensitive.
We represent locks by their concrete addresses, i.e., $\mathsf{Lock}=\mathsf{Address}$, and we define $\mathsf{Lockset}=2^{\mathsf{Lock}}$. Our lockset analysis is implemented as a forward data flow analysis with the domain of locksets, computed for each thread in separation, starting with the empty lockset. On the intraprocedural level, we start to analyse a function with an entry lockset $L_{entry}$ coming from the caller, and gradually update it using the following transfer function that, for a context $c$ with a statement $s$, transforms the lockset $L_{old}$ reached before executing $s$ into a set of locksets $\mathcal{L}_{new}$ resulting from the execution of $s$ as follows:
$\mathcal{L}_{new}=\begin{cases}\{L_{old}\cup\{\ell\}\;|\;\ell\in\mathsf{Backend.value\_ptr}(c,\mathtt{p})\}&\textrm{if s is $\mathtt{lock(p)}$},\\ \{L_{old}\setminus\,\{\ell\}\;|\;\ell\in\mathsf{Backend.value\_ptr}(c,\mathtt{p})\}&\textrm{if s is $\mathtt{unlock(p)}$,}\\ \{L_{old}\}&\textrm{otherwise.}\end{cases}$
After applying the transfer function, if $\mathcal{L}_{\mathrm{new}}$ is not a singleton set, we fork the analysis and continue with each element of $\mathcal{L}_{\mathrm{new}}$ separately. As a result, we obtain the summary of a function $f$ analysed in a context $c$ in the form $(c,L_{entry})\mapsto\mathcal{L}_{exit}$ where $\mathcal{L}_{exit}$ is the set of locksets at the return statements of all forked analysis runs inside $f$. Such summaries are then used for the interprocedural analysis. We also construct statement summaries $\mathcal{S}$ of the same form as for functions, i.e., mappings from pairs of an input context and a lockset to output sets of locksets. These are used for caching at the statement level and to implement the must- and may-lockset queries for a context $c$ as
$\mathsf{may\_ls}(c)=\bigcup_{(c,L)\in\mathsf{dom}(\mathcal{S})}L$ $\mathsf{must\_ls}(c)=\bigcap_{(c,L)\in\mathsf{dom}(\mathcal{S})}L$
@@
As we represent locks using memory addresses and fork the analysis whenever we cannot precisely evaluate the parameter of a (un)locking operation, the set of traversed paths can easily explode. To tackle this problem, our transfer function selects only a small number of locks when the set $\mathsf{Backend.value\_ptr}(c,\mathtt{p})$ is too large (currently, the threshold is set to three locks). This works in many practical cases because it is usually enough to differentiate the may- and must-cases through a few paths only. In the future, we would like to provide support for symbolic treatment of locks combined with may- and must-equality reasoning.
@@
For non-blocking lock functions such as $\mathtt{trylock}$, which does not block when the mutex is already locked, or $\mathtt{timedlock}$, which blocks just for a limited time, we update the lockset with relevant locks for which we also track the variable used to store the return value of the corresponding call. Evaluation of a query $\mathsf{may\_ls}(c)$ or $\mathsf{must\_ls}(c)$ will then ask the backend for the state $s$ computed for the context $c$ through the query $s=\mathsf{Backend.state}(c)$, and filter out the locks for which the state $s$ does not guarantee that the locking was (possibly or surely, depending on the query) successful. This guarantee will only be provided if the success of the locking was explicitly checked in the context $c$. This way we obtain a limited path-sensitivity for branching over expressions containing variables storing return values of lock operations. For the function $\mathtt{pthread\_mutex\_lock}$, we assume that it is blocking (i.e., that it never fails), but the user can easily change the configuration to treat it as non-blocking.
@@
To support read-write locks, a lock object keeps additional information about whether it was obtained in the read- or write-mode. Based on this information, we modify the intersection of locksets, presented later on as the main operation used for data race detection, so that it contains only those locks which are present in both operands and at least in one of them in the write-mode. In other words, if two accesses are guarded only by a lock obtained in both cases in the read-mode, we will treat them as unprotected.
@@ 6.2 Active-Threads Analysis
The sets of active threads are computed in a very similar way to how the lockset analysis is performed – just working with sets of threads instead of sets of locks, with the transfer functions for thread creation and joining analogous to those for locking/unlocking, and with the same way of handling the may- and must-queries. The set of thread entry points is usually small, and we thus do not need any special treatment to tackle path explosion unlike in the lockset analysis. One difference between the analyses is in the initial states. The lockset analysis of each thread is started with the empty lockset, but the active-threads analysis of a thread $t$ starts with the union of all threads that may be active together with $t$ in some context of the parent of $t$ (the analysis is thus run in the order given by a topological order on thread-creation relation, and for the very rare case of cyclic thread-creation graphs, we simply skip the analysis, providing trivial answers for queries based on it).
Thread identifiers.
@@ 7 Data Race Detection
An important part of data race detection is discovering local bases that may escape their scope. Our syntactic backend assumes that dynamically allocated bases can always escape and static bases escape when their address is taken. Using the Alias and EVA backends, we enhance this by testing (for both static and dynamic bases) not only whether their addresses were taken, but also whether they were assigned to global variables.
May- and must-data-races.
We distinguish two types of data races based on the confidence we have in them. For may-data-races, the below conditions need to be satisfied:
  1. 1.
    At least one of the accesses is a write access.
  2. 2.
    The accessing threads are either distinct or the same non-unique thread.
  3. 3.
    The offsets of the addresses may overlap (checked by a query over the $\mathsf{CValue}$ domain of intervals).
  4. 4.
    The intersection of the may-locksets of the concerned contexts is empty (checked using the results of the lockset analysis).
  5. 5.
    The threads of both accesses may run in parallel (checked using the results of the active-threads analysis).
For must-data-races, we require the must-variants of Conditions 3–5 to hold. In addition, the following conditions are required to rule out races involving a memory access representing abstraction of multiple memory accesses:
  1. 6.
    The memory base is not weak (i.e., if it is dynamic, then it represents precisely one dynamic allocation) and it is not allocated in a non-unique thread.
  2. 7.
    The size of the access does not exceed the size of the type of the base.
  3. 8.
    The base is the only base being read/written to in its context.
Condition 6 ensures that we do not report must-races on bases obtained by abstracting and hence merging different calls to allocation functions. Condition 7 rules out races on possibly different indices to arrays (an access with an offset bigger than the size of the type of the base usually abstracts multiple accesses). Finally, Condition 8 handles situations when our context-sensitivity level is not sufficient to distinguish accesses to multiple bases.